OpenStudy (anonymous):

An expression is shown below: 2x3y + 18xy − 10x2y − 90y Part A: Rewrite the expression so that the GCF is factored completely. (4 points) Part B: Rewrite the expression completely factored. Show the steps of your work. (6 points)

OpenStudy (anonymous):

can anyone give me the answer

OpenStudy (anonymous):

@welshfella @Vocaloid

OpenStudy (anonymous):

@Nnesha

OpenStudy (jdoe0001):

hmm what's a GCF?

OpenStudy (anonymous):

greatest common factor

OpenStudy (jdoe0001):

\(\bf 2x^3y+18xy-10x^2y-90y\) <---see any here

OpenStudy (anonymous):

2

OpenStudy (jdoe0001):

so... let us take it out then :) what do you end up with?

OpenStudy (anonymous):

2y(x−5)(x2+9)

OpenStudy (jdoe0001):

well... hmm to be exact.. notice, all terms have a "y" as well, so 2y

OpenStudy (jdoe0001):

hmmm how... did you get 2y(x−5)(x2+9)?

OpenStudy (anonymous):

srry i just used mathway

OpenStudy (jdoe0001):

hmm ok... did the instructions on the exercise said you meant to use mathway and write down the output?

OpenStudy (anonymous):

i just wrote the equation

OpenStudy (anonymous):

an u just ell me the answer for A plz

OpenStudy (anonymous):

really have to get this done

OpenStudy (jdoe0001):

so... take the GCF and see what you end up with :)

OpenStudy (anonymous):

idk

OpenStudy (anonymous):

help plz

OpenStudy (jdoe0001):

\(\bf 2x^3y+18xy-10x^2y-90y\implies {\color{blue}{ 2}}x^3{\color{blue}{ y}}+{\color{blue}{ 2}}\cdot 9x{\color{blue}{ y}}-{\color{blue}{ 2}}\cdot 5x^2{\color{blue}{ y}}-{\color{blue}{ 2}}\cdot 45{\color{blue}{ y}}\) so.. .take out the GCF... and see what's left

OpenStudy (anonymous):

ohhh ok